A house is to be heated by hot water that receives heat from a furnace. Air with an enthalpy of 305 kJ/kg and fuel with an enthalpy of 45360 kJ/kg enters the furnace and the gases leaving have an enthalpy of 608 kJ/kg. There are 17 kg-air per kg-fuel and the water for heating receives 18.3 kW of heat by circulating through the furnace. What is the fuel consumption in kg/hr?


mf = _____ kg/hr

answer in 3 decimal places


1
Expert's answer
2021-08-25T17:24:21-0400

"h_f=45360 \;kJ/kg \\

h_a=305 \;kJ/kg \\

h_g =608 \;kJ/kg \\

Q=18.3 \;kW"

Using conservation of mass:

"m_a+m_f=m_g \\

m_a=21.5m_f \\

21.5m_f+m_f=m_g \\

22.5m_f=m_g"

Energy balance:

"E_{in}=E_{out} \\

m_ah_a+ m_fh_f=m_gh_g + Q \\

21.5m_f \times 305 + m_f \times 45360 = 22.5m_f \times 300 + 18.3 \\

(51917.5 -13680)m_f=18.3 \\

38237.5m_f=18.3 \\

m_f=4.7858 \times 10^{-4} \; kg/sec \times \frac{1}{1/3600 \;hr} \\

m_f= 4.7858 \times 10^{-4} \times 3600 \;kg/hour \\

m_f=1.7229 \;kg/hr"